# Dense SAE Latents Are Features, Not Bugs

### v1

@@ Appendix C Compute Resources Used
We expect the experiments for training SAEs, capturing SAE activations and generating completions with Gemma 2 2B to be able to be run in about 30 A6000 hours. The LLM judging experiments take less than USD \$20 through OpenRouter with Gemini 2.5 Flash Preview [22].
@@ Appendix D Broader Impact
Appendix E Additional Taxonomy Results
E.1 Classification of dense latents
In our taxonomy, we identify dense latents using automated tests. We do not expect these tests to be perfect for a variety of reasons—for instance, dense latents not lining up perfectly with the “true” feature basis due to learning a linear combination basis, and the fundamental difficulty of designing true, causal tests. However, for the purposes of illustration, we choose reasonable cutoffs for each test to create Figure 2, listed below.
  • Context-binding latents: Fraction of successful flips $>0.75$.
  • Nullspace latents: $>0.2$ of encoder weight in bottom 10 $\mathbf{W_{U}}$ singular vector subspace.
  • Meaningful-word latents: AUC of using “is meaningful word” to predict “feature fires” $>0.75$.
  • PC-aligned latents: cosine similarity with top PC $>0.75$.
A few dense latents fall in 2 categories based on our automated tests to find them, with the most common clashes being between sentence- and paragraph- tracking (see § E.2), and between several categories and meaningful-word latent. For the purposes of illustration, we break ties according to the priority (from highest to lowest): {context-tracking, sentence-tracking, alphabet, nullspace, context-binding, paragraph-tracking, meaning, PCA} based on our confidence in our automated tests.
E.2 Position latents
@@
The observation in Figure 15 that period-tracking and newline-tracking latents are hard to distinguish also relates to our discussion in § 6 that because the sparsity incentive is low for these dense latents, they may not be perfectly aligned to “true” model dense features, and may instead be a linear combination of two related features.
E.3 Additional examples of context-binding latents
Figure 16: Normalized run lengths for each latent across layers. Grey line represents the mean. Normalized run length is defined as the average run length, divided by the expected run length under Bernoulli. Context-tracking latents and one latent in layer 24 have been excluded from the plot due to high outlier normalized run lengths.
@@ E.3 Additional examples of context-binding latents
E.4 Steering context-binding latents
Our methodology for steering is as follows:
@@ E.4 Steering context-binding latents
    Prompt Gemma-2-2B with input text from the RedPajama dataset, ending at a natural point (after a newline token), with at least 400 tokens.
@@
  3. 3.
    Allow Gemma-2-2B to generate a completion without steering, and prompt an LLM (Gemini 2.5 Flash Preview) to judge whether the completion is more like F1 activating examples, F2 activating examples, or unclear.
  4. 4.
E.5 Interpreting context-binding latents
We attempt to interpret mid-layer latents that exhibit coherent chunk-level activations, by asking an LLM judge to generate natural language explanations based on activating examples. When examples are drawn from the same context, the explanations are often detailed and specific. However, when examples are drawn from different contexts, the explanations become vague or generic (Table 3).
This drop in specificity across contexts is somewhat expected, since explanations for any SAE feature may overfit the context. It is difficult to rule out the possibility that these dense latents represent an uninterpretable abstract feature the model learns. However, the causal steering experiment seems to cause the relevant specific concepts to be brought up during generation, supporting the “binding” hypothesis.
In-context explanations Out-of-context explanations
References to violence, death, or controversial/negative events, either real or fictional Discussion of a specific, named concept or entity within a broader category (e.g., a specific martial art, a specific plant, a specific architectural movement, a specific medical condition, a specific search tool)
Mentions of genres, particularly speculative fiction genres (sci-fi, fantasy, horror) Lists or enumerations of related items or concepts
References to the Beautiful Creatures book series Identifying or defining a role, title, or specific part of something
References to past events or states of being, often with a negative or challenging connotation Describing a state of being in control, leading, or advancing
Accessing or observing information about others (competitors, other users, etc.) References to specific, named entities or concepts within a larger domain (e.g., a specific TV show, a specific phrase, a specific team name)
References to famous or noteworthy artworks and their associated information (artist, price, era, nickname) Referring to a specific concept or phenomenon by its name or a descriptive phrase
References to online platforms or marketplaces where user-generated content, fan-related items, or resale goods are exchanged References to negative or undesirable actions/events/concepts
Trauma and its impact on mental health, particularly in children Describing a specific curriculum, program, or set of agreements/targets
Mentions of the band Greta Van Fleet or their music Referring to a specific action or event that is happening or has happened
References to totalitarian or authoritarian political ideologies and systems References to specific, identifiable entities or concepts (products, organizations, actions, objects, people)
Table 3: LLM (Gemini 2.5 Flash Preview) generated explanations of suspected context-binding latent L12:7541, given 5 randomly drawn same-context examples vs 5 randomly drawn cross-context examples. 10 example explanations are shown.
E.6 Meaningful-Word Latents
In addition to the high-level categories shown in the main body, we found the AUC for several other categories, some of which are shown here.

### v2

@@ Appendix C Compute Resources Used
We expect the experiments for training SAEs, capturing SAE activations and generating completions with Gemma 2 2B to be able to be run in about 30 A6000 hours. The LLM judging experiments take less than USD \$20 through OpenRouter with Gemini 2.5 Flash Preview [google2025gemini].
@@ Appendix D Broader Impact
Appendix E Additional Taxonomy Results
E.1 Classification of dense latents
In our taxonomy, we identify dense latents using automated tests. We do not expect these tests to be perfect for a variety of reasons—for instance, dense latents not lining up perfectly with the “true” feature basis due to learning a linear combination basis, and the fundamental difficulty of designing true, causal tests. However, for the purposes of illustration, we choose reasonable cutoffs for each test to create Figure˜3, listed below.
  • Context-binding latents: Fraction of successful flips >0.75>0.75.
  • Nullspace latents: >0.2>0.2 of encoder weight in bottom 10 $\mathbf{W_{U}}$ singular vector subspace.
  • Meaningful-word latents: AUC of using “is meaningful word” to predict “feature fires” >0.75>0.75.
  • PC-aligned latents: cosine similarity with top PC >0.75>0.75.
Very few dense latents (3.6% across layers) fall in >1 category based on our automated tests to find them, with the most common clashes being between sentence- and paragraph- tracking (see §˜E.2), and between several categories and meaningful-word latent. For the purposes of illustration, we break ties according to the priority (from highest to lowest): {context-tracking, sentence-tracking, alphabet, nullspace, context-binding, paragraph-tracking, meaning, PCA} based on our confidence in our automated tests.
E.2 Position latents
The observation in Figure˜17 that period-tracking and newline-tracking latents are hard to distinguish also relates to our discussion in §˜6 that because the sparsity incentive is low for these dense latents, they may not be perfectly aligned to “true” model dense features, and may instead be a linear combination of two related features.
@@ E.2 Position latents
E.3 Interpreting context-binding latents
We attempt to interpret mid-layer latents that exhibit coherent chunk-level activations in two ways:
@@ E.3 Interpreting context-binding latents
    All-context: Following existing autointerp methods [paulo2024automaticallyinterpretingmillionsfeatures], we sample 10 activating and 10 non-activating phrases from an entire corpus and ask an LLM (Gemini 2.5 Flash) to generate an explanation. We repeat this 100 times to generate 100 explanations.
@@
    In-context: We instead sample 10 activating and 10 non-activating phrases from the same context. We repeat this 100 times (using 100 different contexts) to generate 100 explanations.
When examples are drawn from the same context, the explanations are specific but highly diverse across contexts. When examples are drawn from different contexts, the explanations become vague or generic (Table˜4). This drop in specificity across contexts is somewhat expected, since explanations for any latent may overfit the context. However, doing the same for sparse latents (Table˜5), we see that a “good” sparse latent would have similar explanations with both in-context and all-context examples, aligning with the usual assumption that SAEs learn directions that represent a concept in the model.
It is difficult to rule out the possibility that these dense latents represent an uninterpretable abstract feature the model learns. However, the steering experiment seems to cause the relevant specific concepts to be brought up during generation, supporting the “binding” hypothesis that there are dense directions that do not represent a fixed concept but rather are used in the model’s computation.
E.4 Additional examples of context-binding latents
We include two additional examples of context-binding latent pairs with high flip score: layer 13 pair (15275, 11449) (Figure˜18) and layer 12 pair (14906, 14599) (Figure˜19). For each pair, we show two example contexts where they are active, illustrating how each latent activates on specific but context-dependent concepts, and that latents in a pair do not co-activate.
@@ E.4 Additional examples of context-binding latents
E.5 Steering context-binding latents
Our methodology for steering is as follows:
  1. 1.
    Prompt Gemma 2 2B with input text from the RedPajama dataset, ending at a natural point (after a newline token), with at least 400 tokens.
  2. 2.
  3. 3.
    Allow Gemma 2 2B to generate a completion without steering, and prompt an LLM (Gemini 2.5 Flash Preview) to judge whether the completion is more like F1 activating examples, F2 activating examples, or unclear.
  4. 4.
E.6 Meaningful-Word Latents
We provide the mapping from higher-level categories to Penn Treebank tags in Table˜3. In addition to the AUC for the categories shown in the text, we show the AUC for a few other categories in Figure˜20, which seem mostly unrelated to predicting the activations of dense latents.
@@ E.7 PCA latents
Figure˜21 reports, layer-by-layer, each dense latent’s cosine similarity to PC1 (left) and the share of its norm contained in the top-5 PCs (right). As noted, only one antipodal pair per layer strongly aligns with PC1; most dense latents place little mass in the top PCs. Figure˜22 shows that this pattern is stable under SAE hyperparameters: varying the sparsity target ($L_{0}$) or dictionary size yields similar PC1 alignments, neither removing nor proliferating PC-aligned latents.
@@
Dense latent L12:7541 ($f=0.421$)
In-context All-context
Numerical values or categories related to real estate listings, such as prices, property types, or number of beds A proper noun or a specific concept/entity within a sentence
Mentions of people, groups, or entities involved in a discussion or action The beginning of a new clause or phrase, often following a comma, parenthesis, or other punctuation, or indicating a shift in topic or focus
Mentions of Donna Brazile or political news and commentary, often critical of establishment figures or media outlets A sequence of words that are part of a proper noun, a title, or a specific phrase, often capitalized, or a short phrase that acts as a label or identifier within a larger text.
References to the Niagara Wire publication or its staff and content The beginning of a new sentence or clause, often following punctuation or a line break
Terms and phrases related to the Cybercrime Prevention Act of 2012 in the Philippines The detection of numbers, dates, or specific numerical references within text
References to Barack Obama or his administration The beginning of a new clause or phrase, often following a conjunction, preposition, or punctuation mark, that introduces additional information or a new element into the sentence structure
Mentions of specific people, places, or things, or references to a particular time or event A proper noun or a common noun that is part of a larger phrase, often appearing after a preposition or as an object of a verb
Mentions of CSIRO’s internal operations, expertise, and collaborative efforts Mentions of specific entities, objects, or concepts within a broader context, often highlighting a particular detail or aspect of the surrounding text
Nautical vessel chartering and related services The letter ’I’ or ’A’ or ’O’ when it is the first letter of a word or a standalone word
Mentions of healthcare organizations, roles, or initiatives related to health and human services A token or sequence of tokens that is part of a larger, multi-word proper noun, compound noun, or specific phrase, where the preceding context helps to complete the meaning of the highlighted part.
Dense latent L12:2009 ($f=0.319$)
In-context All-context
Punctuation marks, numbers, or single letters that are not part of a larger word Code, symbols, or foreign language phrases
Mentions of events, dates, or outcomes related to the life of Stéphanie of Monaco The continuation of a word or phrase across a line break or other formatting boundary
Mentions of the St. John’s Red Storm basketball team, their coach Mike Jarvis, or their player Hatten Short, common words or symbols that are often part of a larger phrase or structure, but do not carry significant meaning on their own
Reporting on COVID-19 cases and related news The beginning of a new sentence or clause, often following a period, comma, or other punctuation, or a line break
Scientific citation formatting and punctuation Punctuation marks, prepositions, and conjunctions that connect different parts of a sentence or list
The beginning of a new clause or sentence The beginning of a new word or token that is not preceded by a space
References to students or pupils in an educational context The detection of a word or phrase that is part of a larger, well-known entity or common expression, where the detected part is not the beginning of the entity or expression
Biographical details and life events of a character, including family, career, and personal status, often presented in a chronological or list-like format Mentions of specific people, places, or entities, or phrases that introduce or refer to them
Specific references to the current or a past UN General Assembly session, or to the UN Secretary General and his staff The beginning of a new sentence or phrase, often following punctuation or a line break, or the start of a new section within a document
Mentions of specific dates, years, or numbers in a historical or official context Mentions of specific words or phrases that are part of a larger list or enumeration, often found in titles, bullet points, or structured content
Table 4: Sampled explanations of dense latents L12:7541 and L12:2009, using in-context examples versus all-context examples. The in-context explanations are highly specific and diverse, while the all-context explanations are vague.
Sparse latent L12:10356 ($f=8.90\times 10^{-4}$)
In-context All-context
Years in the 2010s or 2020s, often following a movie title and sometimes preceded by "HBO" or "HBO Max" A two-digit year following a month or day, or as part of a date range
The number "1" in a four-digit year, specifically in the 2010s decade The last two digits of a four-digit year
The year 2019 in dates or as a standalone year The last two digits of a four-digit year
The digit ’1’ when it is part of a four-digit year that starts with ’20’ and is followed by a two-digit number, typically representing a day or a time, or a forward slash. The second digit of a year in the 21st century
The first digit of a two-digit year in a citation The last two digits of a four-digit year
The first two digits of a four-digit year The last two digits of a year in the 21st century
The release year of a movie title The last two digits of a four-digit year
Mentions of the "Product of the Year" award followed by a specific year The last two digits of a year in a date
The year 2022 in date formats A four-digit year in the 21st century, specifically between 2010 and 2024
The third digit of a four-digit year, specifically when the year is 2013, often found in movie titles or release dates The last two digits of a four-digit year, typically in the 2000s
Sparse latent L12:800 ($f=7.30\times 10^{-4}$)
In-context All-context
Mentions of the South Ossetian conflict, including locations, people, and related events Mentions of people’s names or titles, often followed by their statements or actions
References to a specific person named Sarah, including possessive forms and direct address Mentions of a person or entity speaking or being referenced
Mentions of reports, analyses, or statements made by individuals or groups A proper noun or pronoun that is the subject of a sentence or clause, or a proper noun that is the object of a verb or preposition
Mentions of individuals or organizations involved in mine clearance or humanitarian aid Mentions of people or organizations, often in attribution or as subjects of actions
Mentions of the author Tom Bissell, often in relation to his work or statements Mentions of proper nouns, often names of people or organizations, that are split across a line break
Mentions of "Dicko" as a proper noun, often followed by a verb or punctuation, indicating a new clause or action related to the person. Mentions of people speaking or being quoted
Mentions of Alan Waller, Earl Spencer’s former head of security Proper nouns or pronouns referring to people or organizations, often followed by a verb
Mentions of people or organizations as subjects or possessive entities Mentions of a person’s name followed by a verb of speaking or a reference to that person
Attributions of statements or opinions to individuals or groups, often experts, in news or analytical contexts A proper noun or pronoun that is the subject of a sentence or clause
Mentions of Sonny Dykes, a football coach, or his last name, often in the context of his statements or actions Mentions of people or organizations as subjects or agents of actions
Table 5: Sampled explanations of sparse latents L12:10356 and L2:800, using in-context and all-context examples. While the in-context explanations still tend to be more specific, they still center around a similar theme as the all-context explanations, and it is plausible that L12:10356 is a “date” feature and L12:800 is a “proper noun” feature.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, we claim that we examine the geometry of high frequency latents, taxonomize high frequency latents, and examine their distribution across layers, all of which we do in the paper (we link to these specific sections in the introduction).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have an extensive limitations section in the discussion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We include experimental details sufficient to reproduce the experiments in the main body (all of which use open source models). We provide additional experiment details in Appendix˜B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We plan to upload our code by the supplementary materials deadline.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Justification: As described in the experimental details section, we include experimental details sufficient to reproduce the experiments in the main body (all of which use open source models). We provide additional experiment details in Appendix˜B.
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  30. 7.
    Experiment statistical significance
  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  32. Answer: [No]
  33. Justification: Because we are providing a taxonomy of different SAE latents, statistical significance in most of our experiments does not make sense.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  35. 8.
    Experiments compute resources
  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  37. Answer: [Yes]
  38. Justification: See Appendix˜C for a discussion of the compute used for our experiments.
  39. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code of ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Justification: The research complies in every respect with the code of ethics.
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [Yes]
  48. Justification: See Appendix˜D for a discussion of our paper’s broader impact, which we overall believe to be very positive.
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: We do not release any new data or models.
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: We use and cite the open source Gemma models and GemmaScope SAEs.
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [N/A]
  63. Justification: We do not introduce any new assets.
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and research with human subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: We do not do research with human subjects.
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: We do not do research with human subjects.
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  75. 16.
    Declaration of LLM usage
  76. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  77. Answer: [N/A]
  78. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  79. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.